Problem Solving
12. LC Video rents a game machine for $10 and video games for $3 each. Big Vid rents a game machine for $7 and video games for $4 each.
a) Write a linear equation to represent the total cost of renting a game machine and some video games from LC Video.
b) Write a linear equation to represent the total cost of renting a game machine and some video games from Big Vid.
c) Find the point of intersection of the two lines from part a) and b).
d) Explain what the point of intersection represents in this context.
Thanksssssss and i need it today..

Answer by Mathtut(3670)

12. LC Video rents a game machine for $10 and video games for $3 each. Big Vid rents a game machine for $7 and video games for $4 each.
a) Write a linear equation to represent the total cost of renting a game machine and some video games from LC Video.
b) Write a linear equation to represent the total cost of renting a game machine and some video games from Big Vid.
c) Find the point of intersection of the two lines from part a) and b).
d) Explain what the point of intersection represents in this context.
Thanksssssss and i need it today..
:
a)LC:C(g)=3(g)+10......where g is th number of games rented and C is the total cost.
:
b)BV:C(g)=4(g)+7......where g is th number of games rented and C is the total cost
:
c)C(g)=3(g)+10
..C(g)=4(g)+7
:
set equations equal to each other
:
3g+10=4g+10
:
g=3
:
take g's value and plug it into either equation to find C(g). I choose LC
:
C(g)=3(3)+10
:
C(g)=19
:
so (3,19) are the point of intersection
:
d) this is the point at which both companies have the same cost